# Distributed implementation of tensor-product selected configuration interaction

### v1

@@ 3.1 N2 with the aug-cc-pVDZ Basis Set
480 2259 2484 115 331.2
960 1179 1292 7 344.5
1440 834 917 5 366.8
1920 655 740 14 394.7
2400 554 618 11 412.0
2880 483 552 28 441.6
A particularly notable result is that the total delay time ($T_{\text{delay}}$) does not increase significantly as the number of nodes grows, demonstrating strong parallel efficiency. Generally, increasing the number of nodes worsens communication congestion, which in turn increases MPI latency. In fact, before implementing the optimization strategies introduced in Section 2.4, the measured $T_{\text{delay}}$ values exhibited a significant growth trend with increasing nodes, sometimes even resulting in longer wall times despite using more nodes. However, after applying these strategies, $T_{\text{delay}}$ was significantly reduced, allowing the wall time to decrease consistently with increasing nodes. Although the total node-hours (NHs) still increase with node count (i.e. decreasing parallel efficiency), the substantial wall time reduction confirms the effectiveness of our parallelization strategies.
Table 2: Potential Energy Surface of N2 in Hartree computed with the aug-cc-pVDZ basis set for different bond lengths (R). We report the HF energy and the FCI energies computed with double (FP64) and single (FP32) precision.
$R$ (Bohr) HF FP64 FP32
2.118 -108.956105 -109.297336 -109.297335
2.4 -108.875323 -109.258928 -109.258927
2.7 -108.748038 -109.182160 -109.182160
3.0 -108.618929 -109.108790 -109.108789
3.6 -108.401280 -109.016333 -109.016335
4.2 -108.242916 -109.986038 -109.986044
The FCI energies of the potential energy surface of N2 are shown in Table 2. The double-precision and single-precision results are nearly identical, with the maximum difference being only 0.006 mH at $R=4.2$ Bohr. This data validates the accuracy and reliability of our single-precision implementation for TPSCI calculations.
3.2 CN with the cc-pVTZ Basis Set
As our program is capable of performing open-shell system calculations, we also provide benchmark data for CN, computed with C2v symmetry, frozen core orbitals, and a bond length of 1.16 Å. For the cc-pVDZ and cc-pVTZ basis sets, these calculations require $2.46\times 10^{8}$ and $4.86\times 10^{11}$ determinants, respectively. We emphasize here that naively storing (in single precision) even a single CI vector alone on each node would require about 1.8 TB of memory; however, using our distributed memory scheme, we can unlock the full memory of supercomputer Fugaku and perform such a large calculation. We also report the strong scaling performance of this calculation in Table 3.
Table 3: Parallel performance of TPSCI for CN computed with the cc-pVTZ basis set. We report in seconds the average wall-time (Avg.), maximum wall-time (Max.), and delay time ($T_{\text{delay}}$, see Sec. 2.4.4). We also report the total node hours (NHs) used by the calculation.
@@ 3.2 CN with the cc-pVTZ Basis Set
2400 4019 4400 401 2933
4800 2231 2534 490 3379
7200 1497 1681 242 3362
In Table 4, we present a comparison of the energy values computed by FCI with various orders of coupled cluster theory computed using NWChem 20. The FCI values represent an improvement of approximately 1-2 mH over CCSDT 21, 22, 23. For a system of this size, CCSDTQ 24, 25 provides an excellent approximation, with an error of only about 0.2 mH. In addition, as our calculations represent one of only a few systems that can be computed with a triple-$\zeta$ quality basis set by FCI, we also perform $2/3$ extrapolation of this data to the complete basis set limit (CBS). We extrapolate the HF energy ($E_{c}^{\text{HF}}$) of a given cardinality ($c$) as CBS($l$/$m$) $=E^{\text{HF}}_{m}+\frac{(E^{\text{HF}}_{m}-E^{\text{HF}}_{l})m^{-5}}{m^{-5}-l^{-5}}$ and the correlation energy ($E_{c}^{\text{Cor}}$) as CBS($l$/$m$) $=E^{Cor}_{m}+\frac{(E^{\text{Cor}}_{m}-E^{\text{Cor}}_{l})m^{-3}}{m^{-3}-l^{-3}}$ 26. We hope that this data for a challenging system can be used for benchmarking new methods.
Table 4: Energy values of CN using different methods and basis sets. We denote the two point extrapolation using basis sets of cardinality $l$ and $m$ as CBS($l$/$m$).
Basis Set HF CCSD CCSD(T) CCSDT CCSDTQ FCI
cc-pVDZ -92.197616 -92.475781 -92.489728 -92.490269 -92.491601 -92.491812
cc-pVTZ -92.219490 -92.547372 -92.567450 -92.567478 -92.569054 -92.569252
CBS(2/3) -92.222807 -92.571623 -92.594283 -92.594095 -92.592456 -92.595966
4 Conclusion
@@
Inspired by the idea of TPSCI proposed by an IBM team in 2024, we proposed and implemented a novel and efficient algorithm in this work. To overcome the memory limitations of a single node, we have, for the first time worldwide, implemented distributed storage of the CI vector in an SCI program, allowing the entire memory resources of a supercomputer to be utilized for large-scale computations. Additionally, to further reduce memory usage, we store only the diagonal elements of the Hamiltonian matrix while computing the off-diagonal elements on-the-fly. As a result, we successfully handled up to $4.86\times 10^{11}$ determinants, surpassing the previously reported largest SCI calculation ($2\times 10^{9}$ determinants) by more than two orders of magnitude.
Meanwhile, large-scale parallel computations on supercomputers face a critical challenge: increasing the number of nodes often leads to severe communication congestion, significantly degrading parallel efficiency. To address this, we have developed and implemented a series of optimization strategies tailored to TPSCI, including eliminating unnecessary communication, reducing long-distance data transfers, maintaining balanced computational load and memory distribution in MPI, hiding communication by prefetching, and detecting target process statuses before data transfer. The results demonstrate that even with up to 7,200 nodes, the program’s wall time remains within 116% of the average across all processes, and communication delays ($T_{\text{delay}}$) do not increase significantly with the number of nodes. Such exceptional performance lays a solid foundation for the development of TPSCI in the future.
Since this study is an independent effort of our research group, it does not yet incorporate determinant selection using QC. Therefore, we directly performed FCI calculations on N2 with the aug-cc-pVDZ basis set and CN with the cc-pVTZ basis set, both of which are larger than the cc-pVDZ basis set reported in previous FCI studies of these molecules. These data can be used as benchmark in future quantum chemistry studies.
Looking ahead, the efficient distributed and parallelized framework developed in this work provides a strong foundation for integrating TPSCI with QC approaches. In particular, quantum computers could be leveraged to identify important bit strings, which could then be used to construct a spin-adapted determinant space via tensor products. These determinants would be efficiently processed using our highly efficient and scalable program. The combination of quantum bit string selection with large-scale classical computation has the potential to greatly enhance electronic structure calculations, enabling the accurate simulation of even larger and more complex molecular systems.
@@ Acknowledgments
The authors thank Muneaki Kamiya for suggesting a restart strategy in the Davidson algorithm to reduce memory usage, and Himari Pathak for discussions on applications.

### v2

@@ 3.3 Cr2 with the STO-3G Basis Set
2880 1098 1119 0.4 895
5760 551 567 2.2 907
8640 373 383 7.4 919
11520 281 293 8.1 938
14400 227 242 17.3 968
For large-scale systems, we recommend using the single-precision implementation, as it provides sufficient numerical accuracy while significantly improving MPI efficiency compared to the double-precision variant. Therefore, all subsequent tests focus exclusively on the single-precision variant. As summarized in Table 7, $T_{\text{delay}}$ remains below 18 seconds even when using more than 14,000 nodes. The ratio of maximum to average wall time remains within 107%, and the total NHs increase by less than 9% from 2,880 to 14,400 nodes. These results demonstrate not only near-ideal parallel efficiency but also excellent scalability: even as both the system size and node count grow substantially, our implementation consistently maintains exceptional performance. This underscores the strength and robustness of our MPI optimization strategies in handling large-scale computations.
3.4 N2 with the cc-pVTZ Basis Set
The final system is N2 with the cc-pVTZ basis set at $R=2.068$ Bohr. This calculation involves $2.62\times 10^{12}$ determinants under D2h symmetry with core orbitals frozen, which is approximately twice the size of the largest reported FCI calculation to date ($1.31\times 10^{12}$ determinants).
Table 8: HF and FCI energies (single-precision) of N2 at $R=2.068$ Bohr computed with the cc-pVDZ and cc-pVTZ basis sets. The FCI energies are extrapolated to the complete basis set (CBS) limit using a two-point formula.
cc-pVDZ cc-pVTZ CBS(2/3)
HF FP32 HF FP32 FP32
-108.954553 -109.276527 -108.984093 -109.375153 -109.408722
[…]
Table 9: Parallel performance of TBSCI for N2 computed with the cc-pVTZ basis set using single-precision implementations. We report the average wall time (Avg.), maximum wall time (Max.), and delay time ($T_{\text{delay}}$) of a single distributed matrix-vector multiplication in seconds, as well as the total node hours (NHs) consumed.
@@ 3.4 N2 with the cc-pVTZ Basis Set
12000 2582 2924 0.9 9747
18000 1726 2076 0.3 10380
24000 1346 1630 2.4 10876
30000 1102 1413 12.8 11775
36000 917 1201 8.2 12010
As summarized in Table 9, $T_{\text{delay}}$ remains below 13 seconds even with more than 30,000 nodes, demonstrating the remarkable communication efficiency achieved by our MPI optimization strategies. However, as the number of nodes increases, the ratio of maximum to average wall time becomes significantly larger, reaching 131% at 36,000 nodes. Meanwhile, the total NHs increase about 23% from 12,000 to 36,000 nodes. Considering the use of tens of thousands of nodes, such load imbalance is acceptable, and the overall parallel performance remains excellent. Further improvements to the TBSCI program are planned for future work.
4 Conclusions and Prospects
@@
In this work, we have proposed and implemented a completely new and highly efficient algorithm for the tensor-product bitstring SCI (TBSCI), in which the selected determinants are constructed as tensor products of important bitstrings. 17 To overcome the memory limitations of a single node, we have, for the first time worldwide, implemented distributed storage of the CI vector in the TBSCI program, allowing the entire memory resources of a supercomputer to be utilized for large-scale computations.
To reduce MPI communication congestion, we developed a suite of strategies including avoiding unnecessary communication, leveraging molecular symmetry, reducing long-distance data transfers, balancing computational load and memory usage, overlapping communication and computation through prefetching, reassigning fast-fetch $[0,2]$ tasks, implementing staggered fetch strategies, detecting process status prior to data access, and applying a sleep strategy. These innovations were not developed all at once, but were progressively identified and validated through several months of intensive testing, profiling, and iteration. As these strategies were integrated, the communication delay of the slowest process ($T_{\text{delay}}$) was drastically reduced. The parallel capabilities of TBSCI enabled an SCI-based calculation of N2 using the cc-pVTZ basis set with $2.62\times 10^{12}$ determinants—more than three orders of magnitude beyond the previous SCI record ($2\times 10^{9}$ determinants) 12 and about twice the scale of the largest reported FCI calculation ($1.31\times 10^{12}$ determinants). 5 Despite this scale, the TBSCI single-precision implementation exhibited no communication delay under 10,000 nodes, and maintained an exceptional performance with under 18 seconds delay even at 36,000 nodes, marking a new level of MPI efficiency in large-scale SCI.
Looking forward, a critical next step is the incorporation of second-order perturbative corrections atop the distributed SCI framework. Because TBSCI is the first SCI implementation with distributed CI vector storage, corresponding perturbative corrections must also be implemented in a distributed manner, requiring new algorithms and implementations. This presents substantial technical challenges which will be the subject of future work.
Additionally, we are exploring the development of a fully classical bitstring selection strategy, which would enable TBSCI to operate entirely on classical hardware. Such a version may even outperform hybrid quantum-classical approaches by avoiding the detrimental effects of quantum noise and enabling complete reproducibility.
Finally, the distributed storage scheme and MPI communication optimization strategies developed in this work are not limited to TBSCI. Rather, they constitute a general and scalable framework for large-scale CI implementations, and can be readily adapted to various SCI variants, as well as semistochastic CI methods. 33, 34
We believe this work lays the foundation for a new generation of parallel electronic structure methods capable of addressing strongly correlated systems at unprecedented scale. It also marks a new stage for SCI, driven by distributed CI vector storage and communication optimization strategies.
@@ Acknowledgments
The authors thank Muneaki Kamiya for suggesting a subspace restart strategy in the Davidson algorithm to reduce memory usage, and Himadri Pathak for discussions on applications